# An analysis of nuclear parton distribution function based on Kullback-Leibler divergence

### v1

@@ Abstract
In this work, we present an attempt to quantify the difference between nuclear parton distribution functions (nPDFs) in different nuclei and parton distribution functions (PDFs) in free nucleons using the Kullback-Leibler (KL) divergence, a measure widely employed in quantum information theory. By introducing certain constraints and the “minimum relative entropy” hypothesis, we can determine the shape of the structure function in the intermediate-$x$ region, which is intimately connected with the renowned EMC effect. The obtained results are consistent with the latest global fits of experimental data. Moreover, we have extended this methodology to the determination of the gluon nPDFs, achieving results that exhibit remarkable consistency with global QCD fits. These results indicate that KL divergence may offer us a new window and perspective on the structure and nature of nucleons — particularly in cases where limited experimental and theoretical QCD constraints are available.
@@ I Introduction
Parton distribution functions (PDFs) is a kind of quantum correlation functions which characterize the probability that the parton carries a certain momentum fraction of the hadron momentum. The PDFs serve as fundamental inputs for precision Standard Model (SM) baseline predictions at the LHC and other collider facilities, thus their determination is a key focus in high-energy physics Hou:2019efy ; Bailey:2020ooq ; Lai:2010vv ; Martin:2009iq ; Alekhin:2012ig ; Barone:2000tx .
@@
When free nucleons become bound within atomic nuclei, nuclear parton distribution functions (nPDFs) become necessary. One of the primary goals of nuclear physics is achieving a comprehensive description of the structure of atomic nuclei. The nPDFs, which is responsible for the motion of partons in the nuclear medium, is one of the most poorly understood parts of nuclear structure. A prime example is the unexpected EMC effect, which indicates the quark PDFs in bound nucleons are modified compared with free ones in the intermediate-$x$ region EuropeanMuon:1983wih ; NewMuonNMC:1990xyw ; Gomez:1993ri ; Seely:2009gt . The complex nature of nonperturbative QCD renders the analytical solution of nPDFs unattainable. Furthermore, the lattice QCD also remains impractical on this issue, as nuclear structure represents a quintessential many-body problem that currently exceeds the computational capabilities and resources available for lattice simulations. Even for the very light nucleus, current lattice calculations remain at an exploratory stage Detmold:2020snb ; Amarasinghe:2021lqa ; Chen:2024rgi . Therefore, nearly all current available nPDFs are extracted from global fit analysis of experimental data Eskola:2021nhw ; AbdulKhalek:2022fyi ; Duwentaster:2022kpv ; Helenius:2021tof . Developing new methodologies to pin down nPDFs would be a significant scientific goal.
@@
The recent rapid development of quantum information theory has drawn attention in high-energy physics for its potential applications to old problems, possibly revealing fundamental connections to underlying physical mechanism Galindo:2001ei ; ELEMENTSOFINFORMATIONTHEORY . Quantum information theory has introduced novel observables — various types of entropies that enable an extended characterization of hadron structure or scattering processes beyond classical approaches. Studies in this area have proliferated significantly Wang:2014lua ; Kharzeev:2017qzs ; Beane:2018oxh ; Gotsman:2020bjc ; Fedida:2022izl ; Hentschinski:2022rsa ; Liu:2022grf ; Han:2020vjp ; Ma:2018wtw ; Hagiwara:2017uaz ; Vedral:2002zz ; Nam:2025hei ; Miller:2023ujx ; Zhang:2023oja ; Bloss:2025ywh ; Hu:2025jne ; Wang:2025inr . Here, we need to find an appropriate entropy for the investigation of nPDFs. First, both PDF and nPDF are probability distributions. Second, we want to quantify the difference between these two distributions. The appropriate quantity is the KL divergence $\int p(x)\ln(p(x)/q(x))dx$ (also known as relative entropy) Kullback:1951zyt ; Wehrl:1978zz ; Benito-Calvino:2022kqa ; Cao:2022iqh ; Cao:2022ajt , which is a type of statistical distance: a measure of how much a reference distribution $q(x)$ is different from a true distribution $p(x)$. If we treat the PDF as the reference distribution and the nPDF as the true distribution, their ratio reflects the nuclear modification, whose fundamental mechanism may be recast in the quantum-theoretic terms of KL divergence. When a free nucleon is placed inside an atomic nucleus, its PDF is modified into nPDF due to interactions with surrounding nucleons. Understanding the precise mechanism of this transformation remains a challenge, with ideas like the “minimum relative entropy” hypothesis being explored, as elaborated in this paper below.
In this article, we quantify the difference between PDFs and nPDFs by KL divergence and try to determine the quark structure function in the intermediate-$x$ region. Furthermore, we also applied this method to obtain the less biased gluon nPDF, which is remarkably consistent with the current global QCD fits. Our findings suggest that the KL divergence could serve as a novel and insightful observable for studying nucleon parton structure.
@@
The organization of this paper is as follows. In Sec. II, we give the formalism for structure functions in deeply inelastic scattering (DIS). Sec. III presents the definition of KL divergence and discusses its properties. In Sec. IV, we develop a framework and present results for quantifying and determining quark structure function using the KL divergence. We turn to the investigation of gluon nPDFs in Sec. V. We finish with the conclusion of the paper in Sec. VI.
@@ II Formalism for structure functions in DIS
In experiment, PDFs are extracted from DIS through measurements of structure functions,
$\frac{d^{2}\sigma_{\textrm{DIS}}}{dxdy}=\frac{2\pi\alpha^{2}}{Q^{4}}s\left[1+\left(1-y\right)^{2}\right]F_{2}(x,Q^{2})\,$, (1)
here $x$ is the Bjorken scaling variable, $y$ denotes the inelasticity, $Q^{2}$ represents the negative square of four-momentum transfer from lepton to hadron, and $s$ stands for the square of the center-of-mass energy. The connection between the structure function $F_{2}$ and the quark distributions is given by
@@ III The Kullback-Leibler divergence for structure function
As discussed above, KL divergence is advantageous in analyzing nPDFs from PDFs, it gives us a quantitative measure of the information needed to describe the true distribution (nPDFs) when starting from the reference scenario (PDFs). Having a feeling for the KL divergence, we turn to its physical application on quantifying and determining the nPDFs of quarks and gluons in the next two sections.
IV Framework and results
In this section, we provide calculation method of the KL divergence for different nuclei. The EMC experiment measured the ratio of per-nucleon structure functions for nucleus A to deuteron $(F_{2}^{A}(x,Q^{2})/A)/(F_{2}^{D}(x,Q^{2})/2)$. Here, we take the numerator of this ratio as the true distribution $p^{A}(x)$ and the denominator as the reference distribution $q^{d}(x)$. The $Q^{2}$ dependence is implicitly omitted.
For the reference distribution, deuteron is regarded as an approximation to a free proton and neutron system since it is loosely bound and the average distance between the nucleons is large. Therefore, we adopt the CT18A free PDFs to represent the nucleon nPDFs within the deuteron Hou:2019efy . For the true distribution, we take values from the EPPS21 parametrization for different kinds of nuclei. Since we perform numerical integration of global fitting results, we must renounce to compute the KL divergence across the entire $x$-range $x\in[0,1]$. Here, the range for numerical integration is $x\in[10^{-6},0.99]$, a technical problem arises accordingly is the normalization. As previously introduced, the nonnegativity of KL divergence inherently requires that both distributions be individually normalized to unity. Otherwise, the calculated KL divergence could yield negative values, which is very counterintuitive for a measure representing any sort of entropy. Hence we introduce two normalization factors $z_{p}$ and $z_{q}$, the normalized distributions are
@@ IV Framework and results
The KL divergence between the structure functions of nucleus A and deuteron is calculated thruough

### v2

@@ Abstract
In this work, we propose a method to quantify the difference between nuclear parton distribution functions (nPDFs) in different nuclei and parton distribution functions (PDFs) in free nucleons using the Kullback-Leibler (KL) divergence, a measure widely employed in quantum information theory. By introducing certain constraints and the “minimum relative entropy” hypothesis, we can determine the shape of the structure function in the intermediate-$x$ region, which is intimately connected with the renowned EMC effect. For quark nPDFs, our results align with the latest global fits to experimental data. This agreement suggests that the KL divergence-based methodology may provide novel insight into the structure of nucleons, particularly in cases where experimental data and theoretical QCD constraints are limited, such as those pertinent to gluon nPDFs. Therefore, we applied this methodology to gluon nPDFs, analyzing the results from two commonly used global fitting groups, EPPS21 and nNNPDF3.0. Our analysis reveals that the EPPS21 results align more closely with the “minimum relative entropy” hypothesis. This finding underscores the utility of the proposed method and provides a valuable reference for future global fitting of nPDFs.
@@ I Introduction
Parton distribution functions (PDFs) are a kind of quantum correlation functions which characterize the probability that the parton carries a certain momentum fraction of the hadron momentum. The PDFs serve as fundamental inputs for precision Standard Model (SM) baseline predictions at the LHC and other collider facilities, thus their determination is a key focus in high-energy physics Hou:2019efy ; Bailey:2020ooq ; Lai:2010vv ; Martin:2009iq ; Alekhin:2012ig ; Barone:2000tx .
@@
When free nucleons become bound within atomic nuclei, nuclear parton distribution functions (nPDFs) become necessary. One of the primary goals of nuclear physics is achieving a comprehensive description of the structure of atomic nuclei. The nPDFs, which is responsible for the motion of partons in the nuclear medium, is one of the most poorly understood parts of nuclear structure. A prime example is the unexpected EMC effect, which indicates the quark PDFs in bound nucleons are modified compared with free ones in the intermediate-$x$ region EuropeanMuon:1983wih ; NewMuonNMC:1990xyw ; Gomez:1993ri ; Seely:2009gt . The complex nature of nonperturbative QCD renders the analytical solution of nPDFs unattainable. Furthermore, the lattice QCD also remains impractical on this issue, as nuclear structure represents a quintessential many-body problem that currently exceeds the computational capabilities and resources available for lattice simulations. Even for the very light nucleus, current lattice calculations remain at an exploratory stage Detmold:2020snb ; Amarasinghe:2021lqa ; Chen:2024rgi . Therefore, nearly all current available nPDFs are extracted from global fitting analysis of experimental data Eskola:2021nhw ; AbdulKhalek:2022fyi ; Duwentaster:2022kpv ; Helenius:2021tof .
Moreover, our knowledge of gluon nPDFs is considerably more limited. While modern global fitting analyses yield increasingly precise quark nPDFs—with nuclear modifications for quarks in the intermediate-to-large $x$ region being broadly consistent across different fitting groups—the gluon nPDF remains substantially more uncertain. In contrast to the quark sector, predictions for gluons from different groups often differ markedly. Therefore, developing new methodologies to pin down parton (especially gluon) nPDFs is a significant scientific goal.
@@
The recent rapid development of quantum information theory has drawn attention in high-energy physics for its potential applications to long-standing problems, possibly revealing fundamental connections to underlying physical mechanism Galindo:2001ei ; ELEMENTSOFINFORMATIONTHEORY . Quantum information theory has introduced novel observables—various types of entropies that enable an extended characterization of hadron structure or scattering processes beyond classical approaches. Studies in this area have proliferated significantly Wang:2014lua ; Kharzeev:2017qzs ; Beane:2018oxh ; Gotsman:2020bjc ; Fedida:2022izl ; Hentschinski:2022rsa ; Liu:2022grf ; Han:2020vjp ; Ma:2018wtw ; Hagiwara:2017uaz ; Vedral:2002zz ; Nam:2025hei ; Miller:2023ujx ; Zhang:2023oja ; Bloss:2025ywh ; Hu:2025jne ; Wang:2025inr ; Lin:2025eci . Here, we want to find an appropriate entropy for the investigation of nPDFs. First, both PDF and nPDF are probability distributions. Second, we want to quantify the difference between these two distributions. The appropriate quantity is the KL divergence $\int p(x)\ln(p(x)/q(x))dx$ (also known as relative entropy) Kullback:1951zyt ; Wehrl:1978zz ; Benito-Calvino:2022kqa ; Cao:2022iqh ; Cao:2022ajt , which is a type of statistical distance: a measure of how much a reference distribution $q(x)$ is different from a true distribution $p(x)$. If we treat the PDF as the reference distribution and the nPDF as the true distribution, their ratio reflects the nuclear modification, whose fundamental mechanism may be recast in the quantum-theoretic terms of KL divergence. When a free nucleon is placed inside an atomic nucleus, its PDFs are modified into nPDFs due to interactions with surrounding nucleons. Understanding the precise mechanism of this transformation remains a challenge. This paper explores an information-theoretic approach to this problem, elaborating on how concepts such as the “minimum relative entropy” hypothesis can shed new light on this transformation.
In this paper, we quantify the difference between PDFs and nPDFs by KL divergence and determine the quark nPDFs in the intermediate-$x$ region (EMC region) through minimum relative entropy hypothesis, which proposes that when a free nucleon is bound in a nucleus, its PDFs transform into nPDFs in a way that minimizes the relative entropy between them, subject to known physical constraints. This approach frames the transformation as an optimization problem with fixed boundary conditions, providing a principle to determine shape of nPDFs. In turn, this methodology offers a new benchmark for assessing the quality of global fits to gluon nPDFs, which will be elaborated below.
@@
The organization of this paper is as follows. In Sec. II, we give the formalism for structure functions in deeply inelastic scattering (DIS). Sec. III presents the definition of KL divergence and discusses its properties. In Sec. IV, we develop a framework and present results for quantifying and determining quark structure function using the KL divergence. We turn to the investigation of gluon nPDFs in Sec. V. We finish with the conclusion in Sec. VI.
@@ II Formalism for structure functions in DIS
The cross sections for neutral-current ($eN\to eX$) DIS on unpolarized nucleons can be written in terms of the structure functions
$\frac{d^{2}\sigma_{\textrm{DIS}}}{dxdy}\!=\!\frac{4\pi\alpha^{2}}{Q^{4}}s\Big((1-y)F_{2}(x,Q^{2})+y^{2}xF_{1}(x,Q^{2})\Big)\,$, (1)
here $x$ is the Bjorken scaling variable, $y$ denotes the inelasticity, $Q^{2}$ represents the negative square of four-momentum transfer from lepton to hadron, and $s$ stands for the square of the center-of-mass energy. It should be noted that Eq. (1) only considers the neutral-current photon exchange at leading order (LO) in the strong coupling constant $\alpha_{s}$, and this equation can be further reduced by Callan-Gross relation $F_{2}=2xF_{1}$ Callan:1969uq
$\frac{d^{2}\sigma_{\textrm{DIS}}}{dxdy}=\frac{2\pi\alpha^{2}}{Q^{4}}s\Big(1+(1-y)^{2}\Big)F_{2}(x,Q^{2})\,.$ (2)
This simplified form in Eq. (2) is adopted in this study to facilitate the subsequent analysis using the KL divergence. At LO, the connection between the structure function $F_{2}$ and the quark distributions is given by
@@ III The Kullback-Leibler divergence for structure function
As discussed above, KL divergence is advantageous in analyzing nPDFs from PDFs, and it gives us a quantitative measure of the information needed to describe the true distribution (nPDFs) when starting from the reference scenario (PDFs). Having a feeling for the KL divergence, we turn to its physical application on quantifying and determining the nPDFs of quarks and gluons in the subsequent sections.
IV Framework and results for quark nPDFs
In this section, we provide calculation method of the KL divergence for different nuclei. The EMC experiment measured the ratio of per-nucleon structure functions for nucleus A to deuteron $(F_{2}^{A}(x,Q^{2})/A)/(F_{2}^{d}(x,Q^{2})/2)$. Theoretically, a major advantage of employing this ratio lies in its insensitivity to higher-order perturbative QCD corrections, as the perturbative contributions are largely irrelevant to the nonperturbative PDFs and nPDFs. Here, we take the numerator of this ratio as the true distribution $p^{A}(x)$ and the denominator as the reference distribution $q^{d}(x)$. In this work, $Q^{2}$ is set to $10\,\text{GeV}^{2}$, this is a typical scale in DIS processes. To keep the notation concise, the explicit $Q^{2}$ dependence is implicitly omitted.
For the reference distribution, deuteron is regarded as an approximation to a free proton and neutron system since it is loosely bound and the average distance between the nucleons is large. Therefore, we adopt the CT18A free PDFs to represent the nucleon nPDFs within the deuteron Hou:2019efy . For the true distribution, we take values from the EPPS21 parametrization for different kinds of nuclei  [^1] https://research.hip.fi/qcdtheory/nuclear-pdfs/epps21/ Eskola:2021nhw . Since we perform numerical integration of global fitting results, we must renounce to compute the KL divergence across the entire $x$-range $x\in[0,1]$. Here, the range for numerical integration is $x\in[10^{-6},0.99]$, a technical problem arises accordingly is the normalization. As previously introduced, the nonnegativity of KL divergence inherently requires that both distributions be individually normalized to unity. Otherwise, the calculated KL divergence could yield negative values, which is very counterintuitive for a measure representing any sort of entropy. Hence we introduce two normalization factors $z_{p}$ and $z_{q}$, the normalized distributions are
@@ IV Framework and results for quark nPDFs
The KL divergence between the structure functions of nucleus A and deuteron is calculated through